OpenStudy (yacoub1993):

Find the solutions of the inequality. | x - 5 | ≥ 3

OpenStudy (yacoub1993):

@E.ali @Yttrium

OpenStudy (anonymous):

OK ! The answer off the | x - 5 | is never negative OK ?:)

OpenStudy (yacoub1993):

my question Is why is it never negative?

OpenStudy (anonymous):

Because of | | . If each numbers come in this if they was - we have + and if + have + too ! Just this :)

OpenStudy (yacoub1993):

ok understood let me try to solve it

OpenStudy (yttrium):

 It is (-infinity, 2] U [8,infinity)

OpenStudy (yacoub1993):

how did you do that @Yttrium

OpenStudy (yttrium):

Since it is | | >= 3 You need to create two inequalities x-5 >=3 and x-5 <= -3 And this will lead you to x >= 8 and x <= 2 And then you will form two different lines up to negative and positive infinity. Because one x must be greater than 8 and the other must be less than 2.

OpenStudy (yacoub1993):

ohh ok so how do i write it down on the paper

OpenStudy (yacoub1993):

@Yttrium how do i write it down

OpenStudy (tkhunny):

You may wish to solve the related "less than" problem and remember that you are solving for what is NOT the answer. The Answer: \(|x-5| \ge 3\) NOT The Answer: \(|x-5| \lt 3\) NOT the answer: The Answer: \(-3 \lt x-5 \lt 3\) EXcluded: \(2 \lt x \lt 8\)

OpenStudy (tkhunny):

* second to last one is "NOT the answer" only. Cut and paste problem resulted in that wandering "The Answer".

OpenStudy (yacoub1993):

didn't get what exactly u mean by this "Cut and paste problem resulted in that wandering "The Answer"."

OpenStudy (yacoub1993):

@tkhunny didn't get what exactly u mean by this "Cut and paste problem resulted in that wandering "The Answer"."

OpenStudy (tkhunny):

The third one. It says both "NOT the answer" and "the answer". The second one should not be there.

OpenStudy (yacoub1993):

ok thank you @tkhunny